A business borrowed $2,400 at 7% simple interest per year to buy new tables. If it paid $420 interest, what was the duration of the loan?
Answer by stanbon(75887)

A business borrowed $2,400 at 7% simple interest per year to buy new tables. If it paid $420 interest, what was the duration of the loan?
-----
I = P*r*t
420 = 2400*0.07*t
t = 420/(2400*0.07)
t = 2.5 years
-----------
Cheers,
Stan H.
------------